In a shipment of 10 electronic components, 2 are defective. Suppose that 5
components are selected at random for inspection. What is the probability that:
a. No defective components are included among the 5 selected?
b. Exactly one defective component being included?
1
Expert's answer
2020-11-24T16:46:04-0500

There are 10 components, 2 are defective.

Therefore 10-2=8 components are not defective.

a) Number of ways to choose 5 non defective components is "C(8, 5)"

Number of ways to choose 5 components in total is "C(10, 5)"

Therefore

P(5 non defective components selected) = "\frac{C(8, 5)}{C(10,2)}=\frac{\frac{8!}{5!3!}}{\frac{10!}{5!5!}}=\frac{2}{9}"

b) Exactly one defective component being included: there is one defective and 4 non defective components

Number of ways to choose 4 non defective components is "C(8, 4)"

Number of ways to choose 1 defective component is "C(2, 1)=2"

Number of ways to choose 5 components in total is "C(10, 5)"

Therefore

P(exactly one defective component being included) = "\frac{2\cdot C(8,4)}{C(10,5)}=\frac{\frac{2\cdot8!}{4!4!}}{\frac{10!}{5!5!}}=\frac{5}{9}"


Answer:

a) "\frac{2}{9}"


b) "\frac{5}{9}"